# Mini-Game Lifetime Value Prediction in WeChat

### v1

@@ Abstract.
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 31th ACM SIGKDD Conference on Knowledge Discovery and Data Mining; August 3–7, 2025; Toronto, Canadabooktitle: Proceedings of the 31th ACM SIGKDD Conference on Knowledge Discovery and Data Mining (KDD ’25), August 3–7, 2025, Toronto, Canadaisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Computational advertisingccs: Information systems Collaborative filteringccs: Mathematics of computing Quadratic programming
@@ 3.1. Graph Representation Learning
Obtaining substantial supervisory signals for determining the LTV of individual users is a significant challenge in many practical applications. This difficulty arises from the fact that most users exhibit sparse purchasing behavior, engaging in transactions only infrequently (Chen et al., 2025). One promising approach to mitigate this limitation is the utilization of graph representation learning techniques to capture collaborative signals (He et al., 2020). By leveraging the inherent structure of user-item interaction graphs, these methods enable the learning of robust user and item embeddings even in scenarios characterized by limited availability of direct behavioral data for specific users or items. Recent research has demonstrated the effectiveness of graph representation learning in enhancing the quality of embeddings (Sun et al., 2019). This line of work highlights the potential of graph-based approaches in addressing the challenges posed by sparse user activity data and facilitating more accurate estimation of user LTV.
@@ 3.2. Backbone of GRePO-LTV
The tower layer integrates and processes various embeddings to generate final predictions. It employs AdaSparse (Yang et al., 2022) for domain-specific neuron filtering, adapting the network structure for each domain. To handle the long-tailed distribution of user values, we adopt a zero-inflated lognormal distribution (Wang et al., 2019) modeling approach. We refer our readers to Appendix A for detailed explanations about ZILN.
@@ 3.3. Pareto Optimization
$a$ is the anchoring direction (Mahapatra and Rajan, 2020), and $\boldsymbol{\lambda}\in\mathbb{R}^{m}$ is a predefined weight vector that indicates the importance of each period. Then, we calculate the non-dominating direction $d_{nd}=G\beta^{*}$ and update the model. Therefore, we can yield a Pareto optimal model conditioned on the given importance vector.
Input: Step size $\eta>0$.
1 for $k=1,\ldots,K$ do
2       Randomly generate $u$ and $v$ ranging from $\frac{1}{3}$ to $\frac{2}{3}$;
3       Compute angle $\theta=\frac{\pi}{2}u$, $\phi=\arccos{v}$;
4       Compute weight $\boldsymbol{\lambda}=[\sin{\phi}\cos{\theta},\sin{\phi}\sin{\theta},\cos{\phi}]$;
5       for $t=1,\ldots,T$ do
6             Initialization.
7             Compute gradients of objectives: $G=\nabla\mathcal{L}$;
8             Determine $\boldsymbol{\beta}^{*}$ by solving QP problem in Eq. 1;
9             Calculate non-dominating gradient $d_{nd}=G\boldsymbol{\beta}^{*}$;
10             Update the model with $\eta d_{nd}$;
11       end for
12      Output model $F^{k}$.
13 end for
Output: $F^{*}$ from $[F^{1},\cdots,F^{K}]$.
Algorithm 1 Pareto optimization for multi-horizon LTV
@@ 10             Update the model with $\eta d_{nd}$;
The Pareto optimization for multi-horizon LTV is illustrated in Algorithm 1. The computation of $G^{T}G$ has a time complexity of $O(m^{2}n)$, where $n$ is the dimension of the gradients. Utilizing the most efficient QP solver available (Zhang et al., 2021), the runtime is $O(m^{3})$. Given that in deep networks, it is typically the case that $n\gg m$, the implementation of Pareto optimization does not notably raise the computational expense associated with calculating the non-dominating gradient. Given that the GRePO-LTV components are lightweight, the whole framework remains efficient.
@@ 4.1. Dataset: Wechat Mini-Game Recommendation
Given the absence of public datasets for LTV prediction in a general game platform, we create an industry dataset collected from the WeChat mini games platform to perform offline evaluation. The dataset contains information at three levels: user, game, and behavior. At the user level, we considered each user’s age, gender, city of residence, total number of payments, and other factors. At the game level, we included factors such as game category (e.g., casual), battle type (e.g., level-based), market type (e.g., simulation), and game theme (e.g., music). At the behavior level, we consider the user’s click history for games, which includes detailed information such as the timestamps of clicks and the stay time. To conduct LTV prediction, we retrieve all users who made payments for WeChat mini-games due to advertising over a six-month period and collect their cumulative payments for 3-day, 7-day, and 30-day intervals. Ultimately, we obtain 3,730,392 LTV samples, representing newly registered users for a specific game. On average, there are approximately 21,000 new registered users each day. For offline evaluation purposes, we partitioned the WeChat mini game dataset into training, validation, and test sets with a distribution ratio of 7:2:1.
@@ 4.4. Overall Performance
Our method serves as a strong baseline, consistently outperforming all other baselines by a large margin across the three metrics. Specifically, compared to the second best methods, relative improvements of our method across the three metrics are 14.0%, 3.6% and 1.6%. Overall, models specifically designed for the LTV prediction task tend to outperform classic time series forecasting models. This may be because the former focus on addressing LTV-specific essential challenges such as imbalance, domain shift, sparsity and others.
@@ 6. Related Work
Forecasting the potential revenue of users is critical for online marketing, which directly impacts the effectiveness of advertisements. Early methods rely on simple metrics (e.g., average purchase frequency) and statistical models such as RFM (Fader et al., 2005a) and Pareto/NBD models (Fader et al., 2005b; Bemmaor and Glady, 2012), which focus on historical data and are insufficient to capture complex, non-linear relationships in customer behavior. Machine learning models have been a promising direction to address the above weaknesses and multiple studies have explored them for LTV prediction, including random forest (Vanderveld et al., 2016), XGBoost (Drachen et al., 2018), CNNs (Chen et al., 2018), and RNNs (Xing et al., 2021; Bauer and Jannach, 2021). However, the extreme data sparsity problem has prevented them from achieving better performance, which has gained considerable attention in recent studies. They can be broadly categorized into two classes: (1) leveraging the LTV data from various domains and employing cross-domain learnings to prevent negative transfer (Zhou et al., 2024b; Wang et al., 2024; Weng et al., 2024; Su et al., 2023; Li et al., 2022). For example, CDLtvS (Zhou et al., 2024b) utilizes the LTV data from E-commerce and logistics service domains as well as designs a various-level alignment mechanism to ensure knowledge transfer; (2) Leverage the rich side information to enhance the quality of user representations (Yun et al., 2023; Xing et al., 2021; Yang et al., 2023; Zhang et al., 2023). MDLUR (Yun et al., 2023) employs the additional user portrait and behavior data from multiple sources to enrich the user representations. Moreover, ExpLTV (Zhang et al., 2023) incorporates labels on whether users are highly spent on in-game purchases. In this work, through graph representation learning techniques, we creatively incorporate collaborative signals to obtain better user and game representations.
@@
Pareto optimization has emerged as a powerful tool for modeling trade-offs among multiple objectives in various machine learning tasks, such as neural architecture search (Lomurno et al., 2021) and long-tailed recognition (Zhou et al., 2024c). In recent years, several studies (Jin et al., 2024; Lin et al., 2019; Wu et al., 2022; Xie et al., 2021) have explored its application to multi-objective recommendation systems. PE-LTR (Lin et al., 2019) pioneered the introduction of a Pareto-efficient framework with theoretical guarantees for learning multiple conflicting objectives. Similarly, PAPERec (Xie et al., 2021) proposed a personalized Pareto-efficient framework that jointly predicts the dwell time and click-through rate (CTR) metrics for users. Building upon these advancements, this work leverages Pareto optimization techniques to mitigate the challenges associated with predicting values across different time horizons, thereby enhancing the effectiveness of the recommendation system.
@@ 7. Conclusion
In this paper, we proposed GRePO-LTV, a novel framework for predicting user lifetime value in mini-game advertising. GRePO-LTV addresses data sparsity and multi-period forecasting challenges by employing graph representation learning to capture collaborative signals and Pareto optimization to balance short-term and long-term prediction accuracy. Our method demonstrates significant improvements in offline experiments and online A/B testing. The results highlight the potential of graph representation learning and Pareto optimization in tackling LTV prediction challenges, making GRePO-LTV a valuable tool for optimizing user acquisition, engagement, and monetization strategies in the gaming industry.
Appendix
Appendix A Zero-Inflated Lognormal Distribution
A.1. Overview
@@ A.2. The Lognormal Component
A.3. Structure of the ZILN Distribution
@@ A.4. Mathematical Definition
Appendix B Performance for Different User Groups

### v2

@@ Abstract.
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.2; August 3–7, 2025; Toronto, ON, Canadabooktitle: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.2 (KDD ’25), August 3–7, 2025, Toronto, ON, Canadadoi: 10.1145/3711896.3737248isbn: 979-8-4007-1454-2/2025/08ccs: Information systems Computational advertisingccs: Information systems Collaborative searchccs: Information systems Enterprise applications
@@ 3.1. Graph Representation Learning
Obtaining substantial supervisory signals for determining the LTV of individual users is a significant challenge in many practical applications. This difficulty arises from the fact that most users exhibit sparse purchasing behavior, engaging in transactions only infrequently (Chen et al., 2025). One promising approach to mitigate this limitation is the utilization of graph representation learning techniques to capture collaborative signals (He et al., 2020). By leveraging the inherent structure of user-item interaction graphs, these methods enable the learning of robust user and item embeddings even in scenarios characterized by limited availability of direct behavioral data for specific users or items. Recent research has demonstrated the effectiveness of graph representation learning in enhancing the quality of embeddings (Sun et al., 2019; Li et al., 2025). This line of work highlights the potential of graph-based approaches in addressing the challenges posed by sparse user activity data and facilitating more accurate estimation of user LTV.
@@ 3.2. Backbone of GRePO-LTV
The tower layer integrates and processes various embeddings to generate final predictions. It employs AdaSparse (Yang et al., 2022) for domain-specific neuron filtering, adapting the network structure for each domain. To handle the long-tailed distribution of user values, we adopt a zero-inflated lognormal distribution (Wang et al., 2019) modeling approach. We refer our readers to Appendix B for detailed explanations about ZILN.
@@ 3.3. Pareto Optimization
where $\mathbf{a}$ is the anchoring direction (Mahapatra and Rajan, 2020), $\epsilon>0$ is a hyperparameter, and $\boldsymbol{\lambda}\in\mathbb{R}^{m}$ is a predefined weight vector that indicates the importance of each period. Then, we calculate the non-dominating direction $d_{nd}=G\boldsymbol{\beta}^{*}$ and update the model. Therefore, we can yield a Pareto optimal model conditioned on the given importance vector.
@@
The Pareto optimization for multi-horizon LTV is illustrated in Algorithm A1. The computation of $G^{T}G$ has a time complexity of $O(m^{2}n)$, where $n$ is the dimension of the gradients. Utilizing the most efficient QP solver available (Zhang et al., 2021), the runtime is $O(m^{3})$. Given that in deep networks, it is typically the case that $n\gg m$, the implementation of Pareto optimization does not notably raise the computational expense associated with calculating the non-dominating gradient. Given that the GRePO-LTV components are lightweight, the whole framework remains efficient.
@@ 4.1. Dataset: Wechat Mini-Game Recommendation
Given the absence of public datasets for LTV prediction in a general game platform, we create an industry dataset collected from the WeChat mini games platform to perform offline evaluation. The dataset contains information at three levels: user, game, and behavior. At the user level, we considered each user’s age, gender, city of residence, total number of payments, and other factors. At the game level, we included factors such as game category (e.g., casual), battle type (e.g., level-based), market type (e.g., simulation), and game theme (e.g., music). At the behavior level, we consider the user’s purchase history for games. To conduct LTV prediction, we retrieve all users who made payments for WeChat mini-games due to advertising over a six-month period and collect their cumulative payments for 3-day, 7-day, and 30-day intervals. Ultimately, we obtain 3,730,392 LTV samples, representing newly registered users for a specific game. On average, there are approximately 21,000 new registered users each day. For offline evaluation purposes, we partitioned the WeChat mini game dataset into training, validation, and test sets with a distribution ratio of 7:2:1.
@@ 4.4. Overall Performance
Our method serves as a strong baseline, consistently outperforming all other baselines by a large margin across the three metrics. Specifically, compared to the second best methods, relative improvements of our method across the three metrics are 14.0%, 3.6% and 1.6%. Overall, models specifically designed for the LTV prediction task tend to outperform classic time series forecasting models. This could be attributed to the prior emphasis on resolving fundamental LTV-specific challenges, including imbalance, domain shift, and sparsity.
@@ 6. Related Work
Forecasting the potential revenue of users is critical for online marketing, which directly impacts the effectiveness of advertisements. Early methods rely on simple metrics (e.g., average purchase frequency) and statistical models such as RFM (Fader et al., 2005a) and Pareto/NBD models (Fader et al., 2005b; Bemmaor and Glady, 2012), which focus on historical data and are insufficient to capture complex, non-linear relationships in customer behavior. Machine learning models have been a promising direction to address the above weaknesses and multiple studies have explored them for LTV prediction, including random forest (Vanderveld et al., 2016), XGBoost (Drachen et al., 2018), CNNs (Chen et al., 2018), and RNNs (Xing et al., 2021; Bauer and Jannach, 2021). However, the extreme data sparsity problem has prevented them from achieving better performance, which has gained considerable attention in recent studies. They can be broadly categorized into two classes: (1) leveraging the LTV data from various domains and employing cross-domain learnings to prevent negative transfer (Zhou et al., 2024b; Wang et al., 2024; Weng et al., 2024; Su et al., 2023; Li et al., 2022). For example, CDLtvS (Zhou et al., 2024b) utilizes the LTV data from E-commerce and logistics service domains as well as designs a various-level alignment mechanism to ensure knowledge transfer; (2) Leverage the rich side information to enhance the quality of user representations (Yun et al., 2023; Xing et al., 2021; Yang et al., 2023; Zhang et al., 2023; Liu et al., 2020). MDLUR (Yun et al., 2023) employs the additional user portrait and behavior data from multiple sources to enrich the user representations. Moreover, ExpLTV (Zhang et al., 2023) incorporates labels on whether users are highly spent on in-game purchases. In this work, through graph representation learning techniques, we creatively incorporate collaborative signals to obtain better user and game representations.
@@
Pareto optimization has emerged as a powerful tool for modeling trade-offs among multiple objectives in various machine learning tasks, such as neural architecture search (Lomurno et al., 2021) and long-tailed recognition (Zhou et al., 2024c). In recent years, several studies (Jin et al., 2024; Lin et al., 2019; Wu et al., 2022; Xie et al., 2021) have explored its application to multi-objective recommendation systems. PE-LTR (Lin et al., 2019) pioneered the introduction of a Pareto-efficient framework with theoretical guarantees for learning multiple conflicting objectives. Similarly, PAPERec (Xie et al., 2021) proposed a personalized Pareto-efficient framework that jointly predicts the dwell time and click-through rate (CTR) (Min et al., 2023) metrics for users. Building upon these advancements, this work leverages Pareto optimization techniques to mitigate the challenges associated with predicting values across different time horizons, thereby enhancing the effectiveness of the recommendation system.
@@ 7. Conclusion
We introduced GRePO-LTV, a novel framework for predicting user lifetime value in mini-game advertising. By employing graph representation learning to capture collaborative signals and Pareto optimization to balance short- and long-term accuracy, GRePO-LTV overcomes key challenges like data sparsity and multi-goal conflicts. Our method shows significant improvements in both offline experiments and online A/B testing, highlighting its potential for optimizing game advertising strategies.
Acknowledgements.
This research was supported by Tencent Inc. under project number AMS RBFR2024002. We extend our sincere gratitude to Tencent Inc. for this support.
Appendix A Pareto Optimization for Multi-horizon LTV
Algorithm A1 Pareto optimization for multi-horizon LTV
1:Step size $\eta>0$.
2:for $k=1,\ldots,K$ do
3:     Randomly generate $u$ and $v$ ranging from $\frac{1}{3}$ to $\frac{2}{3}$.
4:     Compute angle $\theta=\frac{\pi}{2}u$, $\phi=\arccos{v}$.
5:     Compute weight $\boldsymbol{\lambda}=[\sin{\phi}\cos{\theta},\sin{\phi}\sin{\theta},\cos{\phi}]$.
6:     for $t=1,\ldots,T$ do
7:         Initialization.
8:         Compute gradients of objectives: $G=\nabla\mathcal{L}$.
9:         Determine $\boldsymbol{\beta}^{*}$ by solving QP problem in Eq. 1.
10:         Calculate non-dominating gradient $d_{nd}=G\boldsymbol{\beta}^{*}$.
11:         Update the model with $\eta d_{nd}$.
12:     end for
13:     Output model $F^{k}$.
14:end for
15:$F^{*}$ from $[F^{1},\cdots,F^{K}]$.
Appendix B Zero-Inflated Lognormal Distribution
B.1. Overview
@@ B.2. The Lognormal Component
B.3. Structure of the ZILN Distribution
@@ B.4. Mathematical Definition
Appendix C Performance for Different User Groups